Find the number of positive integral solution of xyz=330.

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
330 = 2¹ * 3¹ * 5¹ * 11¹
Now
(1+ 3-1) C (3-1) = 3C2 = 3
Hence total number of solutions = 3*3*3*3 = 81
 
Regards
Arun (askIITians forum expert)